elimination method
3x-9y=10
15x-45y=20
I multiplied the first equation by 5 which gives 15x+45y=50
15x-45y=20
___________
30x = 70
is that right?

Answer by tutorcecilia(2152)

Multiply by (-5) instead of (5):
-5(3x-9y=10
15x-45y=20
________________
.
.
-15x+45y=-50
+15x-45y=20
________________
0 = -30 (false statement)
.
[Graphing shows that the lines are parallel, not perpendicular]